Simplify the expression and find its value when a = 5 and b = – 3. 2(a 2 + ab) + 3 – ab

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that a = 5 and b = -3
We have,
= 2a^2+ 2ab + 3–ab
= 2a^2+ ab + 3
Then, substitute the value of a and b in the equation
= (2 × 5^2) + (5 × (-3)) + 3
= (2 × 25) + (-15) + 3
= 50 - 15 + 3
= 53–15
= 38

Thanks